1. I have completed 10000 units of inventory . I would call this 10,000

Beginning work in process

Cost of goods sold

Finished goods inventory

Ending work in process

2. I have sold 10000 units of inventory. I would call this 10,000

Ending work in process

Cost of sold goods

Beginning work in process

Finished work in process

3. I made 10,000 units of inventory. I used direct materials of $60,000, direct labor of $40,000 and had applied factory overhead of $10,000. What was my direct materials cost per unit?

$1

$11

$6

$5

4. If manufacturing overhead is $400,000 and the direct labor hours were 200,000, what is the predetermined overhead rate per labor hour?

$2

50 cents

$4

20 cents

5. The factory employees whose wages would be included in direct labor include

janitors

support personnel

factory supervisors

workers who actually touch and shape the product
Answer #1

1)

The Correct answer is Option (c).

Finished Goods -  Finished Goods means the goods the production of which is completed and which is in hand.

It is said in the question that there is 10000 units in hand production of which is completed. So Option(b) is the correct answer.

(a) Begining Work in Process -

Begining Work In Process means the goods Production of which is started in the last period but not completed within that period. It means the Opening Balance of the Unfinished Goods for this period. But in the Question it is said that the units in hand is finished product and of current period. So, option (a) is incorrect.

(b) Cost Of Goods Sold -

Cost of goods sold means the Cost of the completed Units which is sold.

Example - Suppose a company has 1000 units finished goods in hand the cost of production of which is $10000. From this units 800 units is sold.

So the Cost Of Goods Sold is = (800/1000) X 10000 = $ 8000.

In the question it is said that the units is completed but these units is in hand. Not sold. So, this units are finished goods in hand,not cost of goods sold. Thus option (a) is incorrect.

(d) Closing Work in Process -

Closing Work in process means the goods production of which is started in this year but not completed within current period. In the question the goods are completed. So, Option (d) is incorrect.

(2)

The correct answer is Option (b) Cost of Goods Sold.

In the question it is said that 10000 units is sold. The Cost Related to this 10000 Units is Cost of Goods Sold.

Please refer to Question No.(1) for Cost Of Goods Sold.

(a) Ending Work in Process -

Please refer to Question No.(1) for Ending Work in Process.

It is said in the question that 10000 units are sold. That means 10000 Finished Products are Sold. So, Option (a) is incorrect.

(c) Begining Work in Process -

Please refer to Question (1) for Begining Work in Process.

It is said that 10000 units are sold. that means 10000 finished products are sold. So, Option (c) is incorrect.

(d) Finished Work In Process -

There is no such terms as Finished Work in Process. This option is given just to make confuse.

3) The Correct answer is Option (c)

Cost of Direct Materials Per Unit is $60000/10000 Units = $6 Per Unit

The following options are Incorrect due to the follwing reasons.

Option (a)

This is the Factory Overhead Rate Per Unit. So, this Option is incorrect.

Factory Overhead Rate Per Unit = $10000/10000 units = $ 1/ unit.

Option (b)

This is the Total Cost Per Unit. So, this option is Incorrect.

Total Cost Per Unit = $(60000+40000+10000)/10000 units = $ 11/ unit

Option (d) This option is given to make confusion.

4)

The Correct answer is Option (a).

Predetermined Overhead Rate Per Direct Labour Hour

= Total Factory Overhead / Total DLH = $ 400000 / 200000 DLH = $ 2 Per DLH

From the above calculation it is clearly seen that Option (b), (c) and (d) is incorrect.

5) The Correct Anwer is Option (d).

Direct Labour means the wages paid to the labours who make the product. So, the wages paid to the workers who actually touch and shape the product is to be included in the direct labour.

For the following reasons the below options are not correct

(a) Janitors - Janitors do not make the products. So, wages paid to them should not be included in direct wages.

(b) Support Personnel - Support Personnels don the work for smooth running of the operations of the organisation. They dont make Products. Thus, wages paid to them is not direct labour.

(c) Factory Supervisors - Factory supervisors supervises the factory. They dont make products. So wages paid to them should not be included in the direct labour. So, this option is incorrect.